1. Write an english expression for a. _M_ b. 2a+3
2

2. The area of a rectangular garden is 18m2. If the width is 4.5m, how long is the garden? (A=lw)






Answer by babynarah2007(12)

1. Write an english expression for a. _M_ b. 2a+3
a. _M_ means; Underscore M Underscore

b. 2a + 3 means; Two times a number, increased by three
2. The area of a rectangular garden is 18m2. If the width is 4.5m, how long is the garden? (A=lw)
Rec. Area means: l*w = 18 sqmeters
width is 4.5m means: w = 4.5m
length means, how long is the garden,
so; l*w = 18 sqmeters
l*4.5m = 18 sqmeters
4.5l = 18
l = 18/4.5
l = 4m
Ans, the garden is 4m long
Best luck for your home work... by babynarah2007@yahoo.com